Circular No. 7347

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000B IN NGC 2320 P. Antonini, Avignon, France; and F. Colas and E. Frappa, Pic du Midi

Observatory, report the discovery of an apparent supernova (mag 16.5) on CCD images taken with the Bedouin Observatory 0.305-m telescope by Antonini on Jan. 11.0 UT. SN 2000B is located at α = 7^{h}05^{m}40^{s}.73, δ = +50^{o}35′10′′.5 (equinox 2000.0), which is 14′′ west and 19′′ north of the nucleus of NGC 2320. SN 2000B was confirmed (mag 16.0) via CCD observation on Jan. 13.85 with the 1.05-m telescope at Pic du Midi. The new star does not appear on the Digital Sky Survey (limiting mag ∼ 21) or the Buil-Thouvenot Atlas. Additional photometry of SN 2000B by Colas and Frappa: Jan. 14.03, R = 16.35; 14.80, B = 16.59 ± 0.04; 14.84, V = 16.49 ± 0.02; 14.89, R = 16.26 ± 0.02; 14.94, I = 16.41 ± 0.04; 16.84, V = 16.19; 16.92, R = 16.01; 16.97, I = 16.27; 17.01, R = 16.36. W. Li, University of California, reports that a prediscovery KAIT CCD image taken on Jan. 8.3 shows the new object at mag ∼ 17.5.

PSR 0833–45 R. G. Dodson, P. M. McCulloch, and M. E. Costa, University of Tas-

mania, write: “We observed a sharp glitch on Jan. 16.319 UT in the Vela pulsar. Preliminary model fitting gives a fractional decrease of 3.14 × 10^{−6} in the period, making this the largest decrease on record. The jump was observed at three frequencies (635.034, 990.25, and 1390.64 MHz) with the Mount Pleasant 14-m dish, and further observations were made with the 26-m dish. Details (including temporary parameter files over the glitch and for pre- and post-glitch) can be found at http://kerr.phys.utas.edu.au/

~rdodson/Glitch.”

SUPERNOVAE 1999gh, 1999gk, 1999gm, 1999gn, 1999gq Unfiltered CCD magnitude estimates by L. Kiss, K. Sárneczky, and K.

Sziládi, at the Piszkéstető station of Konkoly Observatory: SN 1999gh in NGC 2986, Jan. 1.020 UT, 16.7 ± 0.3; SN 1999gk in NGC 4653, Jan. 1.181, 16.5 ± 0.2; SN 1999gm in PGC 24106, Jan. 1.039, 18.5 ± 0.3; SN 1999gn in M61, Jan. 1.171, 16.0 ± 0.1; SN 1999gq in NGC 4523, Jan. 1.177, 14.5 ± 0.1.

2000 January 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green